# Content Conditional Debiasing for Fair Text Embedding

### v2

@@ Abstract
Mitigating biases in machine learning models has gained increasing attention in Natural Language Processing (NLP). Yet, only a few studies focus on fair text embeddings, which are crucial yet challenging for real-world applications. In this paper, we propose a novel method for learning fair text embeddings. We achieve fairness while maintaining utility trade-off by ensuring conditional independence between sensitive attributes and text embeddings conditioned on the content. Specifically, we enforce that embeddings of texts with different sensitive attributes but identical content maintain the same distance toward the embedding of their corresponding neutral text. Furthermore, we address the issue of lacking proper training data by using Large Language Models (LLMs) to augment texts into different sensitive groups. Our extensive evaluations demonstrate that our approach effectively improves fairness while preserving the utility of embeddings, representing a pioneering effort in achieving conditional independence for fair text embeddings.
Content Conditional Debiasing for Fair Text Embedding
[…]
@@ 1 Introduction
Embedding text into dense representations is popular in modern Natural Language Processing (NLP), enables applications like sentiment analysis Dang et al. (2020), recommendation systems Zhang et al. (2016), and search engines Palangi et al. (2016). However, as these embeddings become widespread, their inherent biases could affect various applications Packer et al. (2018); Baeza-Yates (2018); Zerveas et al. (2022); Rabelo et al. (2022), potentially harming marginalized groups by perpetuating gender, race, and other biases. Unfortunately, various forms of biases, including gender, racial, and religious biases, have been identified in text embeddings generated by pre-trained language models (PLMs), as reported in several studies Bolukbasi et al. (2016); Nissim et al. (2020); Liang et al. (2020); May et al. (2019). Consequently, attaining fairness in text embedding models is crucial.
@@
Recent debiasing techniques Liang et al. (2020); Kaneko and Bollegala (2021) for text embeddings use post-training to address biases, avoiding the inefficiency of retraining sentence encoders for each new bias. When removing bias, projection-based methods Liang et al. (2020); Kaneko and Bollegala (2021) reduce an embedding’s projection onto each bias subspace. The distance-based method Yang et al. (2023) constructs embeddings for sensitive groups and equalizes distances to text embeddings across these groups. Nevertheless, these methods persist in pursuing independence between sensitive attributes and text embeddings completely remove the sensitive information. As a result, these approaches do not effectively find the sweet spot between fairness and utility trade-off Zhao and Gordon (2022); Deng et al. (2023); Zliobaite (2015).
@@
Recent studies Mary et al. (2019); Deng et al. (2023); Pogodin et al. (2022) suggest that using datasets labeled with sensitive information to achieve conditional independence — specifically, conditioning on the content class to preserve semantic information within the text — provides a more effective approach to achieving fairness while preserving utility. Yet, the scarcity of text datasets with sensitive labels Gallegos et al. (2023) limits the practical application of these findings. Although Counterfactual Data Augmentation (CDA) Zhao et al. (2018) and LLMs Schick and Schütze (2021); Shao et al. (2023) provide ways for generating data, they face challenges with grammatically incorrect artificial instances Dinan et al. (2019) or inherent systematic biases Yu et al. (2023), respectively.
In this paper, we propose a new method to improve text embedding fairness via Content Conditional sensitive attributes Debiasing (dubbed CCD). Specifically, we first use LLMs to augment a dataset into texts across various sensitive groups linked to the same content by substituting sensitive words with equivalents from different groups (i.e., "son", "daughter" and "child" for gender bias). Then, we achieve conditional independence between sensitive attributes and text embeddings by ensuring that texts with varied sensitive attributes but identical content have embeddings equidistant from their neutral counterparts. Extensive evaluations on debiasing benchmarks and downstream tasks demonstrate our method’s effectiveness in promoting fairness while preserving utility.
@@ 2.1 Problem Setting
This section outlines the problem of fairness in text embeddings. We define key variables: $S$ for input text, $C$ for the content, and $f$ for the text embedding model that generates text embedding $C^{\prime}=f(S)$. We also denote sensitive attributes as $A=[a_{1},...,a_{|A|}]$ (e.g., gender, race, age), words related to attribute $a_{i}$ as $X^{a_{i}}$, and neutral words as $X^{n}$. A text is neutral ($S^{n}$) if it lacks sensitive words, and is associated with attribute $a_{i}$ ($S^{a_{i}}$) based on its sensitive polarity Wang et al. (2023). We list all notations in Appendix Table 3.
Fairness Issue: Fig. 1 (a) shows there exists an association between attributes $A$ and content variable $C$, if model $f$ superficially considers $A$ as a proxy for $C$, leading to issues of fairness.
2.2 Content Conditional Fairness Loss
We propose to achieve conditional independence between sensitive attributes and content $A\perp C^{\prime}~{}|~{}C$. The conditional independence allows prediction $C^{\prime}$ to depend on $A$ but only through the content variable $C$, prohibiting abusing $A$ as a proxy for $C$ thus mitigating the fairness issue while preserving the utility. Echoing the methodologies in previous research Hinton and Roweis (2002); Yang et al. (2023), we can define the probability $P(i|j)$ as the likelihood of variable $i$ selecting $j$ as its neighbor in the embedding space. Then conditional independence $A\perp C^{\prime}~{}|~{}C$ can be represented as a content conditioned equal distance on the embedding space, as shown in Fig. 1 (b) and definition 2.1 (detailed proof in Appendix B.5) :
@@ 2.2 Content Conditional Fairness Loss
(Content Conditional Equal Distance Fairness.) Let $S_{C}^{n}$ be a neutral text with the content $C$, assume $S_{C}^{A}=[S_{C}^{a_{1}},S_{C}^{a_{2}},...,S_{C}^{a_{|A|}}]$ being a set of texts from all sensitive groups with the same content $C$. Then embedding model $f$ is content conditioned equal distance fair with respect to attribute $A$, for any $a_{i},a_{j}\in A$:
$||f(S_{C}^{a_{i}})-f(S_{C}^{n})||=||f(S_{C}^{a_{j}})$ $-f(S_{C}^{n})||.$ (1)
To align with this definition of fairness, we design a loss function $L_{bias}$ that aims to mitigate biases while preserving the representation ability of PLMs. For a sample pair $[S_{C}^{a_{1}},...,S_{C}^{a_{|A|}},S_{C}^{n}]$ :
@@
where $dist(A,B)=\exp\left(-\frac{\lVert A-B\rVert^{2}}{2\rho^{2}}\right)$ measures the distance on the manifold Yang et al. (2023); Hinton and Roweis (2002) (details in Appendix B.5), and $\rho$ is selected as the variance of the distance over the training dataset for normalization. To further preserve the valuable information encoded in the model for downstream tasks, we design $L_{rep}$ to enforce similarity between the text embeddings processed by the fine-tuned model $f$ and those processed by the original model $f^{ori}$:
$L_{rep}=||f(S^{n})-f^{ori}(S^{n})||.$ (3)
Thus, the overall training objective is:
$L_{all}=L_{bias}+\beta*L_{rep}.$ (4)
2.3 Content Conditional Data Augmentation
To meet our goal in definition 2.1, we augment a dataset into texts across various sensitive attributes $A$ linked to the same content variable $C$.
Augment Text into Different Sensitive Groups: We facilitate text augmentation by using a pretrained LLM to replace sensitive words with equivalents from different sensitive groups. As shown in Fig. 1 (a), for sensitive words $X^{A}=[X^{a_{i}},...,X^{a_{j}}],a_{i},a_{j}\in A$ in text $S$, the LLM $h$ substitutes $X^{A}$ with words from different sensitive groups $h(X^{A})=[[X_{a_{m}}^{a_{i}},...,X_{a_{m}}^{a_{j}}]]_{a_{m}\in A}$ respectively, creating augmented texts $S^{A}$ from all sensitive groups. We also generate a neutral version $S^{n}$ using neutral replacements, resulting in a dataset of texts from different groups $\mathcal{S}^{A}$ and a neutral text $S^{n}$. Examples are provided in Appendix Table 4.
Rule Guided Prompt Searching: Considering the impact of few-shot prompt tuning on LLM effectiveness, choosing the right prompts is essential for successful text augmentation. As shown in Fig. 1 (d), we first find inaccurately augmented texts whose augmented text’s sensitive group mismatches its sensitive polarity Wang et al. (2023) computed by:
@@ 2.3 Content Conditional Data Augmentation
where $occ$ counts occurrences of words from the sensitive words lists $V^{a_{i}},a_{i}\in A$ in the sentence $S$, we assign polarity based on the most frequently mentioned group in $A$. A text $S^{a_{i}}$ is considered wrongly augmented if $g(S^{a_{i}})\not=i$. We then manually augment such texts to create example prompts for better alignment and repeat the process. This process is iterated to select $K$ (ours 10) samples as few-shot prompts.
3 Experiments
Datasets We generate training data with the News-commentary-v15 corpus Tiedemann (2012). For gender bias evaluation, we follow Yang et al. (2023) to use SEAT May et al. (2019), CrowS-Pairs Nangia et al. (2020) and StereoSet-Intrasentence data Nadeem et al. (2020). We additionally assess fairness on longer texts via the Bias-IR dataset Krieg et al. (2023). To evaluate whether the biased models’ representation ability is maintained, we follow Kaneko and Bollegala (2021); Yang et al. (2023) to select four small-scale subsequent tasks from the GLEU benchmark: Stanford Sentiment Treebank (SST-2 Socher et al. (2013)), Microsoft Research Paraphrase Corpus (MRPC Dolan and Brockett (2005)), Recognizing Textual Entailment (RTE Bentivogli et al. (2009)) and Winograd Schema Challenge (WNLI Levesque et al. (2012)).
Baseline Methods To assess debiasing performance, we compare our algorithm with finetuning-based methods DPCE Kaneko and Bollegala (2021) and ADEPT-F Yang et al. (2023). To assess the effectiveness of our data augmentation strategy, we compare our approach with CDA Zhao et al. (2018). For the use of LLMs in data augmentation, we use ChatGPT and Gemini Team et al. (2023).
Hyperparameters We follow Yang et al. (2023) to conduct experiments on the BERT-LARGE-UNCASED pre-trained model. We use Adam to optimize the objective function. During the debiasing process, our learning rate is 5e-5, batch size is 32, and $\beta$ is 1. Results for DPCE and ADEPT-F are with originally reported hyperparameters Kaneko and Bollegala (2021); Yang et al. (2023). Experiments are conducted on a NVIDIA A100 GPU.
3.1 Results and Analysis
Table 1: Evaluation results on debiasing performance. We test the debiased models on SEAT, CrowS-Pairs, and filtered StereoSet-Intrasentence, with the best result in bold.
[…]
DPCE $-0.21$ $0.27$ $0.44$ $0.07$ $0.25$ $0.21$ 0.24 $81.19$ $56.72$ $65.41$ $64.06$ 52.96 $60.26$ $52.29$
CCD-CDA $0.16$ $0.03$ $0.43$ $0.38$ $0.47$ $0.22$ $0.29$ $80.34$ $\mathbf{53.53}$ $74.69$ $79.10$ $\underline{53.46}$ $73.62$ $46.95$
CCD-GPT $0.47$ $-0.06$ $-0.33$ $0.55$ $-0.17$ $0.48$ $0.34$ $83.76$ $55.40$ $\underline{74.71}$ $82.70$ $54.80$ $\mathbf{74.75}$ $\mathbf{49.62}$
CCD-Gemini $0.47$ $-0.00$ $-0.02$ $-0.72$ $-0.30$ $0.07$ $\underline{0.26}$ $82.91$ $\underline{54.93}$ $\mathbf{74.72}$ $82.97$ $55.00$ $\underline{74.67}$ $\underline{48.85}$
Table 2: Evaluation results on the GLUE dataset and the Bias-IR dataset, we additionally calculate the average deviation to 0.5 for Bias-IR as AVG-DEV. The bold and underline represent the best and second-best respectively.
Datasets Method GLUE $\uparrow$ Bias-IR (Male Ratio, 0.50 the best)
[…]
BERT $92.9$ $84.6$ $\underline{72.5}$ $38.0$ $72.0$ $0.71$ $0.50$ $0.75$ $0.46$ $0.75$ $0.68$ $0.61$ $0.64$ $0.16$
DPCE $92.8$ $69.6$ $53.4$ $49.3$ $66.3$ $0.86$ $0.79$ $1.00$ $0.47$ $0.70$ $0.84$ $0.61$ $0.75$ $0.24$
ADEPT-F $93.2$ $\underline{85.5}$ $69.9$ $56.3$ $76.2$ $0.50$ $0.50$ $0.75$ $0.53$ $0.80$ $0.68$ $0.65$ $0.63$ $0.13$
CCD-CDA $92.8$ 86.3 $65.3$ $50.7$ $73.8$ $0.79$ $0.79$ $0.83$ $0.80$ $0.70$ $0.79$ $0.83$ $0.79$ $0.29$
CCD-GPT 93.6 $85.1$ $70.4$ $\underline{56.3}$ $\underline{76.4}$ $0.50$ $0.43$ $0.58$ $0.60$ $0.35$ $0.53$ $0.61$ 0.51 0.07
CCD-Gemini $\underline{93.5}$ $83.6$ 72.9 56.3 76.6 $0.57$ $0.64$ $0.58$ $0.60$ $0.70$ $0.42$ $0.65$ $\underline{0.59}$ $\underline{0.11}$
Reducing Gender Biases In Table 1, our experiments reveal that our CCD with all data strategies (CDA, GPT, and Gemini) excels in debiasing, consistently outperforming baselines in the StereoSet and the CrowS-Pairs dataset and ranking second best in SEAT. It is worth noting that our method achieves a good ICAT score in the StereoSet dataset, indicating an excellent balance between performance and fairness. However, DPCE, while preserving great fairness, adversely affects its representation capability, as evidenced by a notably lower LMS score in the StereoSet dataset.
Preserving Representation Ability In Table 2, the GLUE results demonstrate that CCD-Gemini achieves the highest average performance, suggesting that it not only maintains but also enhances the model’s representational capabilities. Conversely, DPCE, which is intended to separate gender attributes from neutral text embeddings, adversely affects the model’s utility.
Bias in Information Retrival We conduct an evaluation using the Bias-IR dataset to assess the fairness of text embedding for long text sequences. Table 2 illustrates that our approach, CCD-Gemini and CCD-GPT outperforms others in fairness across numerous fields, as well as in average fairness.
Comparision of Data Strategy To demonstrate the effectiveness of our proposed data strategy, we conduct comparisons with CDA as shown in Table 1 and Table 2. Integrating our debiasing loss with both data strategies resulted in improved fairness. However, CDA experienced a drawback in maintaining model utility, due to potential grammatical inconsistencies introduced by its rule-based augmentation. This highlights the superiority of our LLM-based augmentation method. For different LLMs, we found that both ChatGPT and Gemini achieve good performance on fairness and utility.
Embedding Visualization Fig. 2 reveals that original BERT clusters male and female embeddings, indicating fairness issues. DPCE and ADEPT-F debias by separating sensitive attributes from neutral embeddings, harming the representation quality. Our CCD maintains the original embedding proximity and mixes male and female embeddings, achieving fairness effectively.
4 Conclusion
This paper presents a new method for generating fair text embeddings by ensuring conditional independence between sensitive attributes and text content. We show that conditional independence can be represented as content-conditioned equal distance in embedding spaces. By leveraging LLMs to augment datasets, we address the data scarcity issue and enable model debiasing. Extensive experiments in diverse datasets show that our approach enhances fairness while preserving utility.
5 Limitaions
6 Ethical Consideration
Appendix A Related Work
LLMs in creating dataset: Leveraging the success of large language models (LLMs), researchers have begun using them to generate various forms of training data, such as tabular data Borisov et al. (2022), relation triplets Chia et al. (2022), sentence pairs Schick and Schütze (2021), and instruction data Shao et al. (2023),et al.. As we focus on obtaining data with sensitive attribute information, data generation for text classification would be the most similar one among those applications. Recent efforts in generating data for text classification  Meng et al. (2022); Ye et al. (2022); Wang et al. (2019) primarily employ simple class-conditional prompts while focusing on mitigating issues of low quality after generation. However, these efforts encounter the challenge of inherent systematic biases present in LLMs Yu et al. (2023). Although recent study Yu et al. (2023) considered generated data bias, it focuses only on the diversity of topics and overlooks the inherent bias within words in a text (i.e., "child" occurs more frequently with "mother").
Debias Text Embedding: Bias in text embeddings (also known as sentence embedding) is a significant issue that arises when these models reflect or amplify societal stereotypes and prejudices found in their training data. To resolve the issue,  Liang et al. (2020) contextualizes predefined sets of bias attribute words to sentences and applies a hard-debias algorithm Bolukbasi et al. (2016). Contextualized bebiasing methods Kaneko and Bollegala (2021); Yang et al. (2023) apply token-level debiasing for all tokens in a sentence and can be applied at token- or sentence-levels Kaneko and Bollegala (2021) to debias pretrained contextualized embeddings. However, all the above methods aim to strictly achieve independence between text embedding and sensitive attributes, which may not achieve a good balance between fairness and utility.
Appendix B Algorithm Details
B.1 Notation
Table 3: Main notations used in this paper.
Basic Variables
$L$triangleq
loss function
$f$,$f^{ori}$triangleq
finetuned and original text embedding model.
$h$triangleq
Large language model.
$\theta_{p}$triangleq
Few-shot prompts that used to empower a LLM.
$A$, $a_{i}$triangleq
Sensitive attribute set and $i$-th sensitive attribute.
$S^{a_{i}}$,$S^{n}$triangleq
Text that relate to sensitive attribute $a_{i}$ and neutral text.
$C$,$C^{\prime}$triangleq
Content variable and predicted content.
$X^{a_{i}}$, $X^{n}$triangleq
words from group $a_{i}$ and neutral words in a text.
$V^{a_{i}}$triangleq
words list that contains all collected words related to attribute $a_{i}$.
B.2 Fainess Goals
Improving Fairness There are two goals in improving the fairness of a text embedding model Yang et al. (2023); Kaneko and Bollegala (2021). (1) Remove the superficial connection between the content $C$ and sensitive attributes $A$ so that the fairness is improved (2) Maintain the utility of the text embedding model so that the performance is protected.

### v3

# LLM-Assisted Content Conditional Debiasing for Fair Text Embedding

@@ Abstract
Mitigating biases in machine learning models has become an increasing concern in Natural Language Processing (NLP), particularly in developing fair text embeddings, which are crucial yet challenging for real-world applications like search engines. In response, this paper proposes a novel method for learning fair text embeddings. First, we define a novel content-conditional equal distance (CCED) fairness for text embeddings, ensuring content-conditional independence between sensitive attributes and text embeddings. Building on CCED, we introduce a content-conditional debiasing (CCD) loss to ensure that embeddings of texts with different sensitive attributes but identical content maintain the same distance from the embedding of their corresponding neutral text. Additionally, we tackle the issue of insufficient training data by using Large Language Models (LLMs) with instructions to fairly augment texts into different sensitive groups. Our extensive evaluations show that our approach effectively enhances fairness while maintaining the utility of embeddings. Furthermore, our augmented dataset, combined with the CCED metric, serves as an new benchmark for evaluating fairness.
LLM-Assisted Content Conditional Debiasing for Fair Text Embedding
Wenlong Deng1,2thanks: Work done at google., Blair Chen2, Beidi Zhao1, Chiyu Zhang1, Xiaoxiao Li1thanks: Equal Corresponding Author., Christos Thrampoulidis1† Department of Electrical and Computer Engineering1, The University of British Columbia, Vancouver, BC, Canada {dwenlong,beidiz,chiyuzh,xiaoxiao.li,cthrampo}@ece.ubc.ca Google Cloud2, USA {chenblair}@google.com
@@ 1 Introduction
Embedding text into dense representations is a widely used technique in modern NLP, powering applications such as sentiment analysis Dang et al. (2020), recommendation systems Zhang et al. (2016), and search engines Palangi et al. (2016). However, the extensive use of these embeddings introduces inherent biases that can affect various applications Packer et al. (2018); Baeza-Yates (2018); Zerveas et al. (2022); Rabelo et al. (2022). For instance, search engines Huang et al. (2020) preprocess all text contents and search queries into embeddings to optimize storage and enable efficient similarity matching. These inherent biases in text embeddings can influence the calculation of embedding similarity, impacting the filtering of numerous documents to find pertinent ones. Moreover, text embeddings are directly employed in other applications such as zero-shot classification Yin et al. (2019); Radford et al. (2021) and clustering John et al. (2023). Unfortunately, various forms of biases, including gender, racial, and religious biases, have been identified in text embeddings generated by pre-trained language models (PLMs), as reported in several studies Bolukbasi et al. (2016); Nissim et al. (2020); Liang et al. (2020); May et al. (2019). Consequently, attaining fairness in text embedding models is crucial.
@@
Recent debiasing techniques Liang et al. (2020); Kaneko and Bollegala (2021) for text embeddings use post-training to address biases, avoiding the inefficiency of retraining sentence encoders for each new bias. When removing bias, projection-based methods Liang et al. (2020); Kaneko and Bollegala (2021) reduce an embedding’s projection onto each bias subspace. The distance-based method Yang et al. (2023) constructs embeddings for sensitive groups and equalizes distances to text embeddings across these groups. Nevertheless, these methods persist in pursuing independence between sensitive attributes and text embeddings, which results in the complete removal of sensitive information. As a result, these approaches do not effectively find the sweet spot between fairness and utility trade-off Zhao and Gordon (2022); Deng et al. (2023); Zliobaite (2015).
@@
Recent studies Mary et al. (2019); Deng et al. (2023); Pogodin et al. (2022) suggest that using datasets labeled with sensitive information to achieve conditional independence — specifically, conditioning on the content class to preserve semantic information within the text — provides a more effective approach to achieving fairness while preserving utility. Yet, the scarcity of text datasets with sensitive labels Gallegos et al. (2023) limits the practical application of these findings. To create such datasets, Counterfactual Data Augmentation (CDA) Zhao et al. (2018) collects sensitive-related words and employs a rule-based method to augment the data, but this approach encounters challenges due to the need for an extensive list of words. Finally, while Large Language Models (LLMs) Schick and Schütze (2021); Shao et al. (2023) have offered new methods for data generation thanks to their rich contextual knowledge, yet they still struggle with inherent systematic biases Yu et al. (2023).
In this paper, we improve the text embeding fairness through defining fairness with theoretical analysis, a novel debiasing loss design, and an LLM-based data strategy for dataset generation. Our contributions include:
  • Introducing CCED fairness for text embeddings, ensuring equal sensitive information and conditional independence between sensitive attributes and embeddings.
  • Proposing CCD loss to achieve the desired CCED fairness by ensuring that texts with varied sensitive attributes but identical content have embeddings equidistant from their neutral counterparts.
  • Employing LLMs to augment datasets fairly, representing diverse sensitive groups within the same content for effective training with CCD. Proposing polarity-guided prompting to ensure the LLM-generated data quality and minimize the potential biases from LLMs.
  • Establishing CCED fairness as a benchmark for evaluating fairness in text embeddings.
  • Extensive evaluations on debiasing benchmarks and downstream tasks demonstrate CCD’s effectiveness in promoting fairness while preserving utility.
2 Related Work
Debias Text Embedding: Bias in text embeddings (also known as sentence embedding) is a significant issue that arises when these models reflect or amplify societal stereotypes and prejudices found in their training data. To resolve the issue,  Liang et al. (2020) contextualizes predefined sets of bias attribute words to sentences and applies a hard-debias algorithm Bolukbasi et al. (2016). Contextualized debiasing methods Kaneko and Bollegala (2021); Yang et al. (2023) apply token-level debiasing for all tokens in a sentence and can be applied at token- or sentence-levels Kaneko and Bollegala (2021) to debias pretrained contextualized embeddings. However, all the above methods aim to strictly achieve independence between text embedding and sensitive attributes, which may not balance fairness and utility well. While Shen et al. (2021, 2022) employ contrastive learning losses to mitigate biases in language representations for text classification, their approach relies on supervised data, which is often scarce and expensive to obtain, and primarily focuses on fairness in the subsequent task. Additionally, although Leteno et al. (2023); Shen et al. (2022) observe that representational fairness and group fairness in subsequent tasks are either not correlated or only partially correlated, it is important to note that fairness in subsequent tasks and fairness in text embeddings are distinct areas, with the latter being crucial for various applications. A detailed discussion of these differences can be found in Appendix A.2. In this paper, we utilize LLMs to augment training data for learning fair text embeddings with proposed CCD loss.
LLMs for Dataset Generation: Leveraging the success of LLMs, researchers have begun using them to generate various forms of training data, such as tabular data Borisov et al. (2022), relation triplets Chia et al. (2022), sentence pairs Schick and Schütze (2021); Zhang et al. (2024), and instruction data Shao et al. (2023); Wu et al. (2024). As we focus on obtaining data with sensitive attribute information, data generation for text classification would be the most similar one among those applications. Recent efforts in generating data for text classification  Meng et al. (2022); Ye et al. (2022); Wang et al. (2019) primarily employ simple class-conditional prompts while focusing on mitigating issues of low quality after generation. However, these efforts encounter the challenge of inherent systematic biases present in LLMs Yu et al. (2023). While Yu et al. (2023) considers generated data bias, it focuses only on the diversity of topics and overlooks the inherent bias within words in a text (e.g. ‘child’ occurs more frequently with‘mother’). In this paper, we instructs the LLM to only locate the gendered words and replace them with counterparts from other groups and propose polarity-guided prompt searching to minimize biases from LLMs and ensure the quality of augmented data.
@@ 3.1 Problem Setting
This section outlines the problem of fairness in text embeddings. We define several key variables: $S\in\mathcal{D}$ represents the input text from the data distribution, $C$ denotes the content of the text, [^1] For instance, the texts ‘he is a teacher’ and ‘she is a teacher’ both convey the same content $C$ = ‘is a teacher’. and $A=[a_{1},\ldots,a_{|A|}]$ represents the sensitive attributes (e.g. gender and age). The symbol $n$ indicates neutral, meaning no sensitive information is present. A text with content $C$ is considered neutral $S_{C}^{n}$ if it contain no sensitive information, whereas text $S_{C}^{a_{i}}$ is associated with the sensitive attribute $a_{i}$ if its sensitive polarity Wang et al. (2023) is $a_{i}$, see Eq. (6). The text embedding model $f$ processes a text into a $d$-dimensional embedding $Z\in\mathbb{R}^{d}$. The embedding of a neutral text encodes the content information $C^{\prime}$ (a well trained model $C^{\prime}\approx C$), while the embedding of a sensitive text additionally encodes sensitive information. Words in the text related to the attribute $a_{i}$ are denoted as $X^{a_{i}}$, and neutral words are denoted as $X^{n}$. For clarity, we provide detailed notations in Table 8 in Appendix.
Fairness Issue: Fig. 1 (a) shows there exists an association between attributes $A$ and content variable $C$. If model $f$ superficially treats $A$ as a proxy for $C$ [^2] For instance, raising children is frequently associated with women in the training corpus, resulting in the proxy effect., it results in encoded $C^{\prime}$ being represented by $A$ thus embedding $Z$ will mainly contain sensitive information, which leads to issues of fairness.
Fairness Goal: Mitigating fairness is not trivial, as we need to address not only bias mitigation but also the protection of the model’s representation ability. As shown in Fig. 1 (a), our method aims to (1) break the association between content $C$ and the sensitive attribute $A$, and (2) preserve useful sensitive information in the text embedding. For example, in the case of a text about a father raising a child, its embedding should retain information about the father.
3.2 Content Conditional Debiasing
To break the superficial association, we propose to achieve conditional independence between sensitive attributes and content $A\perp C^{\prime}~{}|~{}C$. The conditional independence allows prediction $C^{\prime}$ to depend on $A$ but only through the content variable $C$, prohibiting abusing $A$ as a proxy for $C$ thus mitigating the fairness issue while preserving the utility. To protect utility, our objective is not to completely remove sensitive information but to ensure that text embeddings from different sensitive groups with identical content contain an equal amount of sensitive information.
3.2.1 Fairness Definition
Firstly, we propose a novel content conditional equal distance fairness for fair text embedding:
@@ 3.2.1 Fairness Definition
(Content Conditional Equal Distance (CCED) Fairness.) Let $S_{C}^{n}$ be a neutral text with content $C$. Assume $S_{C}^{A}=[S_{C}^{a_{1}},S_{C}^{a_{2}},...,S_{C}^{a_{|A|}}]$ being a set of texts from all sensitive groups with the same content $C$. Then, embedding model $f$ is content conditioned equal distance fair with respect to attributes $A$, for any $a_{i},a_{j}\in A$:
$\|f(S_{C}^{a_{i}})-f(S_{C}^{n})\|=\|f(S_{C}^{a_{j}})$ $-f(S_{C}^{n})\|,$ (1)
[…]
As shown in Fig. 1 (b), CCED fairness requires that texts with the same context from different sensitive groups have equal distance to their corresponding neutral text on the embedding space. This text embedding fairness definition has two merits:
Equal sensitive information: The equal distance to the neutral embedding ensures an equitable encoding of sensitive information across diverse groups, allowing fair usage of sensitive information and preserving the utility of embeddings.
Content Conditional Independent: Echoing the methodologies in previous research Hinton and Roweis (2002); Yang et al. (2023), the conditional independence $A\perp C^{\prime}~{}|~{}C$ can be represented as the CCED on the embedding space:
Assumption 3.2.
(Equal Probability) Within a content $C$, the likelihood $P(a_{i}|C)$ on all sensitive attributes $a_{i}\in A$ is uniform $P(a_{1}|C)=...=P(a_{A}|C)$.
Theorem 3.3.
When the equal probability assumption holds, achieving content conditioned equal distance fairness is equivalent to achieving conditional independence between sensitive attributes and content $A\perp C^{\prime}~{}|~{}C$.
Assumption 3.2 is true for a fair dataset that has balanced texts from all groups within content $C$ (can be obtained through our data augmentation strategy in Section 3.3). Theorem 3.3 demonstrates the merit of CCED fairness (Definition 3.1) in achieving embedding fairness. Detailed proof can be found in Appendix A.5.
3.2.2 Content Conditional Debiasing Loss
Based on the defined CCED fairness, we design a loss function $L_{bias}$ that aims to mitigate biases while preserving the representation ability of PLMs. For a sample pair $[S_{C}^{a_{1}},...,S_{C}^{a_{|A|}},S_{C}^{n}]$ :
@@ 3.2.2 Content Conditional Debiasing Loss
where $dist(A,B)=\exp\left(-\frac{\lVert A-B\rVert^{2}}{2\rho^{2}}\right)$ measures the distance on the embedding manifold Yang et al. (2023); Hinton and Roweis (2002) (details in Appendix A.5), and $\rho$ is selected as the variance of the distance over the training dataset for normalization. To further preserve the valuable information encoded in the model and achieve efficient debiasing, we design $L_{rep}$ to enforce high similarity between the neutral texts’ embeddings processed by the fine-tuned model $f$ and those processed by the original model $f^{org}$:
$L_{rep}=\|f(S^{n})-f^{org}(S^{n})\|.$ (3)
Ensuring that neutral embeddings remain unchanged offers two benefits: preserving the model’s representational capability and maintaining neutral embeddings as a consistent reference point in the debiasing loss, ensuring stable equal distance to embeddings with various sensitive attributes. Thus, the overall training objective is:
$L_{all}=L_{bias}+\beta*L_{rep},$ (4)
where $\beta$ is a hyper-parameter used to balance the two terms. An ablation study for setting $\beta$ is detailed in Table 7.
3.3 LLM-Assisted Content Conditional Data Augmentation
We leverage the rich contextual knowledge of LLM with few-shot prompting to obtain a dataset that (1) fulfills the Assumption 3.2 to achieve our goal in Definition 3.1 as well as (2) avoids introducing inherent bias in LLM to augmented data. The data augmentation algorithm is shown in Alg. 1, followed by a detailed explanation below.
Algorithm 1 Data Augmentation Algorithm
Input: Dataset $\mathcal{D}$, Sensitive word lists $V$, Pretrained LLM $h$, Task Description $T$, Example Prompts $P$.
1:for $k$ in $1,\ldots,K$ do $\triangleright$ $K=10$ in this work
2:     Block I: Augment Texts into Different Sensitive Groups
3:     for $S\in\mathcal{D}$ do
4:         $h(S,T,P)\rightarrow[S^{a_{1}},...,S^{a_{|A|}},S^{n}],c$
5:     end for
6:     if $k=K$ then
7:         return Augmented Dataset $\mathcal{D}^{\prime}$
8:     end if
9:     Block II: Polarity Guided Prompt Searching
10:     for $[S^{a_{1}},...,S^{a_{|A|}},S^{n}]\in D^{\prime}$ do
11:         Polarity Checking Eq.6
12:     end for
13:     Manually Augment the wrong augmentation with highest $c$ and add to $P$.
14:end for
Augment Text into Different Sensitive Groups: As shown in Fig. 1 (c), our task description $T$ instructs the LLM to only locate the gendered words and replace them with counterparts from other groups, leaving the other content unchanged thus avoiding fairness issues in text generation. Specifically, for sensitive words $X^{A}=[X^{a_{i}},...,X^{a_{j}}],a_{i},a_{j}\in A$ in the text $S$, the LLM $h$ substitutes $X^{A}$ with words from different sensitive groups and neutral terms, thus obtaining augmented texts from all sensitive groups (as shown in Table 1):
$h(S,T,P)=[S^{a_{1}},...,S^{a_{|A|}},S^{n}],c$ (5)
where $c$ is the confidence score and $P$ is the example prompts (detailed in Table 10 in Appendix). After augmentation, the dataset will have an equal amount of texts from each sensitive group with identical content, meeting our equal probability Assumption 3.2.
Polarity-Guided Prompt Searching: To ensure the quality of augmented texts and the effectiveness of few-shot prompt tuning on LLMs, finding appropriate prompts $P$ is crucial. We propose identifying difficult samples from incorrectly augmented texts to use as prompts. First, these incorrectly augmented samples are detected through a sensitive polarity check as described by Wang et al. (2023) and illustrated in Fig. 1(d). By counting the occurrences of words in predefined sensitive word lists $V=[V^{a_{i}},...,V^{a_{j}}],a_{i},a_{j}\in A$, the polarities of a series of sentences are determined as follows:
@@ 3.3 LLM-Assisted Content Conditional Data Augmentation
where $occ$ represents the number of times words from the list $V^{a_{i}}$ appear in all augmented sentences $S$. For a properly augmented sentence $S^{a_{i}}$, its polarity should match the sensitive attribute $a_{i}$. If $g(S^{a_{i}})\neq a_{i}$, the sentence is considered inaccurately augmented. Then we introduce our prompt searching strategy in Algorithm 1. In each iteration, the algorithm identifies the incorrectly augmented sample with the highest confidence $c$, manually augments it, and adds it to the example prompts $P$. This rule-guided prompt search is repeated $K$ times (with $K=10$) to prepare samples for the few-shot prompt tuning of de-biasing LLMs.